Shutdown activities

Sub detectors

Timetable
   <<June >>
Mo
17
Tu
18
We
19
Th
20
Fr
21
07:00
08:00
09:00
10:00
11:00
12:00
13:00
14:00
15:00
16:00
17:00
18:00
19:00
20:00
     
07:00
08:00
09:00
10:00
11:00
12:00
13:00
14:00
15:00
16:00
17:00
18:00
19:00
20:00

  • 09 Oct 2009 - 09 Oct 2009 - 15:00-24:00 Closing of the cavern for the TED run (UX85B)
  • 12 Oct 2009 - 13 Oct 2009 - 00:00-24:00 TED run, cavern closed (UX85B)
  • 19 Oct 2009 - 19 Oct 2009 - 15:00-24:00 DSO test, cavern closed (UX85B)
  • 20 Oct 2009 - 21 Oct 2009 - 00:00-24:00 DSO test, cavern closed (UX85B)
  • 23 Oct 2009 - 23 Oct 2009 - 15:00-24:00 Closing of the cavern for Injection test (UX85B)
  • 24 Oct 2009 - 26 Oct 2009 - 00:00-24:00 Closing of the cavern for Injection test (UX85B)
  • 07 Oct 2009 - 08 Oct 2009 - 08:00-24:00 RICH HPD exchange, side A
  • 09 Oct 2009 - 09 Oct 2009 - 08:00-15:00 RICH HPD exchange, side A, Survey
  • 14 Oct 2009 - 14 Oct 2009 - 14:00-18:00 RICH HPD exchange, side C
  • 15 Oct 2009 - 18 Oct 2009 - 08:00-18:00 RICH HPD exchange, side C
  • 19 Oct 2009 - 19 Oct 2009 - 08:00-15:00 RICH HPD exchange, side C, Survey

Sub detector Task Responsible start_DATE end_DATE Reference COMMENTS STATUS
  template       TEMPLATE   roll eyes (sarcastic) in preparation
Magnet checking position of coils on supports R. Lindner 12 January 2009 16 January 2009 MX0001   big grin done
Beam Pipe Beam pipe protection removal/installation for M1 survey P. Lepeule 08 January 2009 09 September 2009 VX0001 Protection back on 9.01 at 17:00 big grin done
RICH 1 Photon Funnel, dismount C. Frei 20 January 2009 30 January 2009 RU0001   big grin done
RICH 1 REMOVAL OF THE RICH1 UPPER HPD ARRAY B. Cameron 09 January 2009 15 January 2009 RU0002 Handling big grin done
Tracker Turicensis Work on electronics boxes and opening of TT station for film crew J. van Tilburg 09 February 2009 13 February 2009 TT0001 ongoing big grin done
Calorimeter Change of Turbine HCAL side A R. Kristic 13 January 2009 13 January 2009 CE0001   big grin done
Infrastructure Scaffolding for RMS measurements B. Corajod 12 January 2009 14 January 2009 IX0001 scaffolding big grin done
Muon PERPARATION FOR/AND INSTALLATION OF CHAMBER SUPPORTS, MODIFICATION TO FLEXIBLE GAS PIPES B. Schmidt, A. Cardini 13 January 2009 20 February 2009 UX0001 yellow resp. white nacelle big grin done
Muon FINALIZATION OF AIR COOLING STATIONS M2-M5 B. Schmdt 13 January 2009 20 February 2009 UX0002   big grin done
RICH 2 Check of the gas tighness of the gas enclosure of the RCH2 C. Frei 06 January 2009 26 February 2009 RD0001   big grin done
RICH 1 Installation of the Lower Photon Funnel on the RICH1 detector C. Frei 02 February 2009 06 February 2009 RU0003   big grin done
Outer Tracker Installation of a new access platform close to the magnet which will replace the existing one C. Fournier 12 January 2009 06 March 2009 TO0001 green nacelle big grin done
Inner Tracker TEST SERVICE BOXES ON IT-1C AND IT-3C F. Blanc 16 January 2009 20 January 2009 TI0001   big grin done
VELO Work on the cooling system close to the VELO E. Jans 05 January 2009 16 January 2009 TV0001 scaffolding big grin done
Calorimeter Access to HCAL/C to commission CW/PMT Iouri Guz 16 January 2009 20 January 2009 CE0002 yello nacelle big grin done
Outer Tracker INSTALL CTRLBOX PLATFORMS C-SIDE A. Pellegrino 26 January 2009 30 January 2009 TO0002   big grin done
Inner Tracker SERVICE BOX TILTING MECHANISM (IT-2A) F. Blanc 20 January 2009 21 January 2009 TI0002   big grin done
Inner Tracker SERVICE BOXES ON IT-1A,2A,3A F. Blanc 23 January 2009 27 January 2009 TI0003   big grin done
Calorimeter ECAL INSTALLATION OF HV-BOXES ON C-SIDE Irina Machikhiliyan 20 January 2009 23 January 2009 CE0003   big grin done
Calorimeter Access to HCAL/C to commission CW/PMT Iouri Guz 21 January 2009 26 January 2009 CE0004   big grin done
Calorimeter Check of the PS VFE electronics P.Perret 22 January 2009 06 February 2009 CE0005   big grin done
Calorimeter The installation of PMT/C-W I. Machikhiliyan 26 January 2009 06 February 2009 CE0006 yellow nacelle big grin done
Calorimeter Work on Calorimeter front-end electronics Frédéric Machefert 22 January 2009 22 January 2009 CE0007 Work on Calo platform big grin done
Infrastructure Bunker extension access on side C and A C. Fournier 26 January 2009 14 February 2009 IX0002 Access on bunker extension restricted big grin done
Calorimeter TUNING OF LED SYSTEM Evgeny Gushchin 21 January 2009 09 February 2009 CE0008 Green nacelle, ECAL platform & RCH 2 tower big grin done
Calorimeter CALO SPD VFE 2 REPAIR E. P. Olloqui 27 January 2009 30 January 2009 CE0009 M1 closed on C-side big grin done
Outer Tracker INSTALL HEATING BLANKETS ON T2-Q02-VX A. Pellegrino 03 February 2009 20 February 2009 TO0003 heating T2-Q02-VX for about 10 days big grin done
Calorimeter COMMISSIONING OF ECAL C-SIDE I. Machikhiliyan 02 February 2009 06 February 2009 CE0010 Yellow Nacelle big grin done
HCAL swap 2 cables of CW/PMT in the central part Iouri Guz 06 February 2009 06 February 2009 HC0003 White nacelle, side A big grin done
PS Repair of a broken pin in a PS backplane crate A-side. Patrick Robbe 05 February 2009 05 February 2009 CP0005   big grin done
RICH 1 Installation of trolley with the HPD array B. Cameron 11 February 2009 12 February 2009 RU0004   big grin done
RICH 1 Installation of the Upper Photon Funnel Christoph Frei 16 February 2009 25 February 2009 RU0005 Green Nacelle big grin done
PS Tuning of LED light pulse hight on PS Evgeny GUSHCHIN 10 February 2009 17 February 2009 CP0006 Green nacelle big grin done
ECAL INSTALLATION OF HV-BOXES ON A-SIDE Irina Machikhiliyan 16 February 2009 20 February 2009 CE0011   big grin done
Muon PERPARATION/MODIFICATION OF CALBING FOR CHAMBERS OF REGION R1 AND R2 B. Schmidt, A. Cardini 26 January 2009 24 Mars 2009 UX0003   big grin done
Muon INSTALLATION OF LONG ALIGNMENT BARS FOR STATION M1 B. Schmidt, Th. Schneider 17 February 2009 02 Mars 2009 UX0004 yellow nacelle big grin done
Muon M1 CHAMBER INSTALLATION B. Schmidt, A. Cardini 16 February 2009 20 June 2009 PPSPS-M1-Chamber-Installation http://lhcb-tech-coor.web.cern.ch/lhcb-tech-coor/Task_description/M1_closing.wmv big grin done
RICH 1 HPD array to the detector C-side B. Cameron 25 February 2009 26 February 2009 RU0006   big grin done
ECAL Opening A-side; Beginning of CALO access A-side R. Kristic 26 February 2009 02 March 2009 CE0012 green nacelle in center big grin done
Inner Tracker Repair electronics on IT-2A F. Blanc 25 February 2009 27 February 2009 TI0004   big grin done
Calorimeter The installation of PMT/C-W Irina Machikhiliyan 27 February 2009 13 March 2009 CE0013   big grin done
PS Check of the VFE electronics P.Perret 02 March 2009 13 March 2009 CP0007   big grin done
SPD Cable extension for a control clock cable of the SPD Eduardo Picatoste Olloqui 26 February 2009 27 February 2009 CP0008   big grin done
RICH 2 Installation of the LAMS Didier Piedigrossi 27 February 2009 6 March 2009 RD0002   big grin done
Infrastructure Dismount 2nd level of RICH 2 tower side C C. Fournier 9 March 2009 11 March 2009 IX0003   big grin done
PS/SPD TUNING OF PS/SPD LED SYSTEM E. GUSHCHIN 26 February 2009 15 March 2009 CP0009   big grin done
Inner Tracker Install Service Box tilting F Blanc 11 March 2009 27 March 2009 TI0005   big grin done
Infrastructure Dismount top level of RICH 2 tower side A C. Fournier 16 March 2009 18 March 2009 IX0004 During operation, OT/IT/M1/PS/SPD closed big grin done
RICH 2 Replacement of HPD side C C. Frei 18 March 2009 25 March 2009 RD0003 M1 & Calo closed on that side big grin done
RICH 2 Replacement of HPD side A C. Frei 25 March 2009 1 April 2009 RD0003 M1 & Calo closed on that side big grin done
VELO Maintenance of the three turbo pumps E. Jans 24 March 2009 26 March 2009 TV0002   big grin done
SPD CALO BROKEN CONTROL/CLOCK CABLE Eduardo Picatoste Olloqui 17 March 2009 20 March 2009 CP0010   big grin done
RICH 1 OPENING THE GAS ENCLOSURE FOR SURVEY AND REPAIRS. B. Cameron 26 March 2009 27 March 2009 RU0007   big grin done
ECAL ECAL HV CONTROL BOARD REPLACEMENT A. Konoplyannikov 25 March 2009 25 March 2009 CE0014   big grin done
Muon Replacement of some chambers of Muon stations M2-M5 B. Schmidt 23 March 2009 30 April 2009 UX0005   smile started
ECAL DISMOUNTING OF INTEGRATORS, ECAL C-SIDE I. Machikhiliyan 6 April 2009 20 April 2009 CE00015 2h access big grin done
Inner Tracker Repair VCSELs on digitizer boards F. Blanc 15 April 2009 24 April 2009     big grin done
Calorimeter Closing of CALO A-side for M1 R. Kristic 22 April 2009 22 April 2009 CE00016 at 13:00 big grin done
Tracker Turicensis Exchange TT modules and work on service boxes Jeroen van Tilburg 29 June 2015 03 July 2009 TT0002   roll eyes (sarcastic) in preparation
MOEDAL Preparation work D. Lacarrere, R. Soluk 13 May 2009 15 May 2009 DD0001   big grin done
Infrastructure Installation of access under balcony C. Fournier 28 April 2009 29 May 2009 IX0005   big grin done
HCAL REPLACEMENT RADSOURCE CASSETTE R. Dzhelyadin 22 May 2009 22 May 2009 HC0004 ~ 3h big grin done
Magnet EPDM insertion between upstream coil support and coil C. Fournier 25 May 2009 29 May 2009 MX0002   big grin done
Infrastructure ABSOLUTE GRAVITY MEASUREMENTS S. Guillaume 29 June 2009 30 June 2009 IX0006 UX85A, close to PZ big grin done
SPD CALO SPD CONTROL CABLE REPARATION P. Perret 3 June 2009 4 June 2009 CP0011 needed vertical nacelle big grin done
ECAL Opening ECAL A-side Robert Kristic 19 June 2009 19 June 2009 CE0017   big grin done
ECAL Install modified integrators on ECAL A-side Irina Machikhiliyan 19 June 2009 22 June 2009 CE0018   big grin done
PS Change of the PS VFE LV patch panels Pascal PERRET 26 June 2009 6 July 2009 CP0012   big grin done
ECAL INSTALLATION OF INTEGRATORS, ECAL C-SIDE Irina Machikhiliyan 2 July 2009 6 July 2009 CE0019   big grin done
VELO - Changing pressure and temperature sensors Eddy Jans 6 July 2009 10 July 2009 TV0003   big grin done
Infrastructure High Fog OT phase 1 E. Thomas 10 August 2009 14 August 2009 IX004 No work between magnet and RICH2 big grin done
Inner Tracker Repair cables in RMS patch panel F. Blanc 6 August 2009 13 August 2009 IT0007   big grin done
RICH 2 CLOSURE OF THE GAS ENCLOSURE C. Frei 14 August 2009 14 August 2009     big grin done
ECAL Replacement of malfunctioning PMT/C-W doublets Irina Machikhiliyan 2009-08-18 2009-08-19 CE0020   big grin done
MOEDAL Put up a few test panels and drill holes on top of VELO Richard SOLUK 2009-09-14 2009-09-16 DD0002   big grin done
ECAL Replacement of one fuse in LV chain (A-side) and two cable connectors (C and A side) Irina Machikhiliyan 2009-09-04 2009-09-04 CE0021   big grin done
ECAL MAINTENANCE OF ECAL VOLTAGE DISTRIBUTION/LED TRIGGER SYSTEMS Irina Machikhiliyan 2009-09-29 2009-09-30 CE0022   roll eyes (sarcastic) in preparation
RICH 2 Replacement of HPD side A C. Frei 2009-10-07 2009-10-09   M1 & Calo closed on that side. roll eyes (sarcastic) in preparation
RICH 2 Replacement of HPD side C C. Frei 2009-10-14 2009-10-19   M1 & Calo closed on that side. roll eyes (sarcastic) in preparation
ECAL Repair the ECAL HV_control Anatoly Konoplyannikov 2009-10-05 2009-10-05 CE0023   big grin done

-- RolfLindner - 05 Jan 2009

Topic attachments
I Attachment History Action Size Date Who Comment
Microsoft Word filedoc CE0015.doc r1 manage 44.0 K 2009-04-06 - 20:05 RolfLindner  
Unknown file formatdocx High-Fog-OT-work_statement.docx r1 manage 37.0 K 2009-08-11 - 08:47 EricThomas High Fog installation over OT bridge
Microsoft Word filedoc IT0007.doc r1 manage 108.5 K 2009-08-07 - 13:57 EricThomas Inner Tracker work week 32-33
Microsoft Word filedoc PPSPS-M1-ChamberInstallation.doc r1 manage 232.5 K 2009-02-17 - 09:36 BurkhardSchmidt PPSPS M1 Chamber Installation
Microsoft Word filedoc UX0003.doc r1 manage 111.5 K 2009-02-17 - 09:55 BurkhardSchmidt UX0003
Microsoft Word filedoc UX0004.doc r1 manage 111.0 K 2009-02-17 - 10:03 BurkhardSchmidt  
Microsoft Word filedoc UX0005.doc r1 manage 110.0 K 2009-03-30 - 20:30 BurkhardSchmidt  
Edit | Attach | Watch | Print version | History: r138 < r137 < r136 < r135 < r134 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r138 - 2011-05-27 - RolfLindner
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LHCb All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright &© 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback